Government spending is 700; investment is 310 and consumption is given by 

C=250+0.8Yd

Net tax T=-50+0.25Y

I. Cal equilibrium income

Ii. Assume investment falls by 80units,compute new level of equil income and the value of budget deficit/surplus at that level.of income.

Iii. Suppose that government raises lump-sum taxes by the any of deficit /surplus in order to bal the dudget,what will be the new level.of equil income?


1
Expert's answer
2021-02-26T10:35:02-0500

C= 250+0.8yd

Yd= Y-T

= Y-(-50+0.25y)

= Y+50-0.25y

Therefore, C= 250+0.8(y+50-0.25y)

C= 290+0.6y , G= 700 and I= 310

Where C = consumption function

G= Government spending and I= investment.

Let Y represent nation income

Therefore the equilibrium national=C+I+G

I.e Y= C+I+G

Y= 290+0.6y+310+700

Y-0.6y= 290+310+700

Y(1-0.6) = 1300

Y= 1300 divided by 0.4

Y= 3250.


ii. If investment falls by 80 units, therefore new investment (I) = 310-80=230.

Therefore, new national income is determined as thus;

Y= 290+0.6y+230+700

Y-0.6y= 290+230+700

Y(1-0.6) = 1220

Y= 1220 divided by 0.4

Y= 3050.

The new national income output is 3050.


Government revenue= -50+(0.25 multiply by 3050)

Government revenue= 712.5

Budget surplus= 712.5 - 700 = 12.5.


III. Tax function when lump sum is raised by 12.5 surplus is determined as thus;

Recall that,T= -50+0.25y

Therefore, new Tax function (T) = -50+12.5+0.25y

= -37+0.25y.

New consumption function is determined below:

C= 250+0.8(-37+0.2y)

C= 220+0.2y


New National equilibrium income is determined as thus;

Y= 220+0.2y+230+700

Y-0.2y = 220+230+700

Y(1-0.2)= 1150

Y= 1150 divided by 0.8

Y= 1437.5.